Jack was 17 years older than Pracilla. Together their ages totaled 159 years. What was their age?
Answer by macston(5194)

P=Priscilla's age; J=Jacks age=P+17
P+J=159 Substitute for J
P+(P+17)=159 Subtract 17 from each side.
2P=142 Divide each side by 2
P=71 ANSWER 1: Priscilla is 71 years old.
J=P+17=71+17=88 ANSWER 2: Jack is 88 years old.